OpenStudy (anonymous):

Write the equation in slope- intercept form. What are the slope and y-intercept? -9x+10y=-9 My answer was y=9/10x-9/10; slope: 9/10; y-intercept: -9/10

OpenStudy (anonymous):

y=mx+b m=slope b=y intercept

OpenStudy (anonymous):

in a nutshell, everything looks in place

OpenStudy (anonymous):

does that mean its correct though? haha

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cause m=slope=9/10 and b=y intercept= -9/10

OpenStudy (anonymous):

and your equation is correct as well

OpenStudy (anonymous):

thank you @jacobrocket173

OpenStudy (anonymous):

np @jadareese